OpenStudy (anonymous):

Express the complex number in trigonometric form. 6 - 6i

OpenStudy (anonymous):

\[a-bi= A \cos q\] avec \[A=\sqrt{(a^{2}+b ^{2)}} \] et \[q=\tan^{-1}(-b/a) \]

terenzreignz (terenzreignz):

Je ne parle pas pas francais :) ... Anyway, Let's start with a+bi, instead... @suivpasyonw is quite correct with how to find the value of A. But... for q, which I'll call the argument, it is a little trickier than that, it's safe to first take \[\large r = \tan^{-1}\frac{|b|}{|a|}\] And then, to find q, there are 4 cases... I'm going to use degrees here, okay? If a and b are BOTH POSITIVE q = r. If a is NEGATIVE and b is POSITIVE q = 180 - r If a and b are BOTH NEGATIVE q = 180 + r If a is POSITIVE and b is NEGATIVE q = 360-r ca va? XD

OpenStudy (anonymous):

I am so sorry @terenzreignz. I speak english as a second language and just forgot to write my answer in english :-)

OpenStudy (anonymous):

So the answer would be \[6\sqrt{2}(\cos7\pi/4+isin7\pi/4)\]

OpenStudy (anonymous):

I wish @jazzie u got my answer!

OpenStudy (anonymous):

Yay !! Thank you (:

terenzreignz (terenzreignz):

Good on you :)

terenzreignz (terenzreignz):

@suivpasyonw There were only two French words, and you used mostly Maths symbols, I don't think there was really a need to apologise :D

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

Good @jazzie